OpenStudy (hkmiu):

x^2 - y^2 = 4 in polar coordinates

hartnn (hartnn):

to convert cartesian to polar plug in \(x = r\cos \theta \\ y = r \sin \theta \)

hartnn (hartnn):

and remember that \(\cos^2 \theta -\sin^2 \theta = \cos 2\theta \)